# Hamiltonian stochastic Lie systems and applications

### v2

@@ Abstract
This paper provides a practical approach to stochastic Lie systems, i.e. stochastic differential equations whose general solutions can be written as a function depending only on a generic family of particular solutions and some constants related to initial conditions. We correct the stochastic Lie theorem characterising stochastic Lie systems proving that, contrary to previous claims, it retains its classical form in the Stratonovich approach. Meanwhile, we show that the form of stochastic Lie systems may significantly differ from the classical one in the Itô formalism. New generalisations of stochastic Lie systems, like the so-called stochastic foliated Lie systems, are devised. Subsequently, we focus on stochastic Lie systems that are Hamiltonian systems relative to different geometric structures. Special attention is paid to the symplectic case. We study their stability properties and devise the basics of a stochastic energy-momentum method. A stochastic Poisson coalgebra method is developed to derive superposition rules for Hamiltonian stochastic Lie systems. Potential applications of our results are shown for biological stochastic models, stochastic oscillators, stochastic Lotka–Volterra systems, Palomba–Goodwin models, among others. Our findings complement previous approaches by using stochastic differential equations instead of deterministic equations designed to grasp some of the features of models of stochastic nature.
Keywords: energy-momentum method, Hamiltonian stochastic Lie system, Poisson coalgebra, stability, stochastic Lie system, superposition rule, Stratonovich formalism.
@@ 1 Introduction
In its most classical definition, a superposition rule is a $t$-independent function that describes the general solution of a $t$-dependent system of first-order ordinary differential equations (ODEs) in normal form, a so-called Lie system, via a generic family of its particular solutions and a set of constants related to initial conditions [9, 17, 18, 56]. Superposition rules are used, for instance, in approximate and numerical methods, as they are available for Lie systems whose exact general solutions in explicit form are unknown [20, 51, 56]. In particular, the knowledge of a particular finite set of exact and/or approximate solutions of a Lie system permits one to study its general solution via superposition rules [56].
@@
The superposition rule concept traces its origins back to Sophus Lie’s pioneering and celebrated book [44], redacted and edited by Georg Scheffers. In that work, Lie stated the nowadays called Lie theorem characterising Lie systems. Previous to that, Lie briefly described his Lie theorem in [43] without a proof as a criticism to some previous works on superposition rules by Vessiot, Guldberg, Köningsberger, and other researchers (see [23, 34, 54] and references therein). Lie stated that the results on superposition rules for differential equations on $\mathbb{R}$ devised by previous authors were a simple application of his theory on infinitesimal transformation groups. Moreover, [44] laid down the foundations for the theory of Lie systems. The Lie theorem is also called Lie–Scheffers theorem, as Scheffers took part in writing Lie’s work [44], or Lie superposition theorem [10]. Previous remarks suggest one to say ‘Lie theorem’ rather than ‘Lie–Scheffers theorem’ and to use the denomination ‘Lie system’ instead of ‘Lie–Scheffers system’. Since Vessiot devised important contributions to the theory of Lie systems [20, 23], Lie–Vessiot system is also a good denomination for a Lie system. On the contrary, Scheffers never worked on Lie systems by himself [23].
@@
The Lie theorem states that a $t$-dependent system of ODEs in normal form admits a superposition rule if and only if it describes the integral curves of a $t$-dependent family of vector fields that can be considered as a curve in a finite-dimensional Lie algebra of vector fields, a Vessiot–Guldberg Lie algebra (see [10, 17, 18] for modern approaches and further details).
Lie systems have been thoroughly studied due to their widespread occurrence in physics and mathematics (see [20, 23] counting more than 300 references on Lie systems and related topics). Towards the 1980s, Winternitz and his colleagues at the Centre de Recherches Mathématiques of the University of Montreal conducted an extensive study on Lie systems. Subsequently, many other authors followed their work [23]. There has been a bast effort to generalise Lie systems to more general situations: $t$-dependent Schrödinger equations [19, 20], quasi–Lie systems [16] and quasi–Lie families [14], foliated Lie systems [17, 21], discrete differential equations [50], stochastic differential equations [39], superdifferential equations [8], et cetera [13]. In this work, we are mainly concerned with the extension of superposition rules and Lie systems to the realm of stochastic differential equations [39].
@@
Stochastic differential equations may describe phenomena that deterministic differential equations can not [2]. For example, the probability of disease extinction or outbreak, the quasi-stationary probability distribution, the final size distribution, and the expected duration of an epidemic are features that cannot effectively be modelled by deterministic methods [2]. These and other reasons motivate the huge interest in studying stochastic differential equations. Nevertheless, some deterministic differential equations can grasp certain interesting characteristics of models even easier than stochastic differential equations [12] and stochastic differential equations offer an alternative view that can be complementary [33].
@@
It is interesting to extend superposition rules to stochastic differential equations. The work [39] extends the superposition rule notion to a class of systems of stochastic first-order ordinary differential equations, called stochastic Lie systems. The authors use a Stratonovich approach, as this makes the approach similar to the deterministic theory of Lie systems. Moreover, [39] presents an interesting and very precise presentation of certain local results about stochastic Lie systems and, as a byproduct, it also explains many technical results on standard Lie systems, which use to be absent in the literature [17, 18, 23]. Despite its mathematical interest, many of the technical details given in [39] are frequently omitted as they have, in general, not many practical applications. As noted in [10, 23], standard works on Lie systems, even theoretically ones, are essentially interested in local aspects and generic points, which makes them skip many technical details analysed in [39]. Notwithstanding, there is mathematical interest in global and local technical aspects, as illustrated by the work [10] for global superposition rules and [39]. It is worth stressing that stochastic Lie systems were found to have applications in the description of Brownian motions, economical models like the Black–Scholes theory of derivatives prices, and so on [39]. The potential interest of stochastic Lie systems in epidemic models was very briefly mentioned in [27], without, as far as we know, any further development.
@@
In [39], a stochastic Lie theorem characterising Stratonovich stochastic first-order ordinary differential equations admitting a superposition rule was devised, but it contains a mistake that changes its meaning and potential applications. More exactly, the direct part of the stochastic Lie theorem in [39] states that a stochastic Lie system may admit, locally, a superposition rule if its associated Stratonovich operator is related to a family of vector fields spanning an involutive distribution. Our present work proves that the vector fields must additionally span a finite-dimensional real Lie algebra of vector fields, which is a much stronger condition appearing already in the classical Lie–Scheffers theorem [17, 18, 44]. We also determine the precise point of the mistake in [39], we correct the statement of the stochastic Lie theorem, and we provide a counterexample, based on SIS epidemic models, to illustrate when the stochastic Lie theorem in [39] fails. It is worth stressing that the differences between the wrong and the correct versions of the stochastic Lie theorem have no impact in the applications carefully studied in [39].
@@
Moreover, our work presents a concise introduction to stochastic Lie systems, aiming at providing a practical approach while avoiding technical details that are not necessary for general purposes. In this sense, it simplifies the elegant and mostly rigorous mathematical treatment in [39] by using standard assumptions in mathematical constructions. For instance, we focus on local results at generic points, which significantly simplifies previous needed techniques.
@@
Our correction of the stochastic Lie theorem shows that the stochastic Lie theorem has no exclusive stochastic features in the Stratonovich approach: it retains the conditions of the classical Lie theorem. Meanwhile, it is stressed that the conditions for a stochastic differential equation in the Itô form [53] to become a stochastic Lie system do not follow the standard form expected from the deterministic Lie theorem [17, 18, 20]. This is very important practically, as many relevant stochastic differential equations are given in the Itô form and they have to be translated into the Stratonovich approach [52] to apply the methods of our work and [39]. In this respect, the relation between the Itô and the Stratonovich approaches is reviewed and some examples with potential applications are provided in this work. It is worth noting that stochastic differential equations in the Itô form may seem to be stochastic Lie systems, but they are not. This occurs in certain SIS epidemiological models [32].
@@
Apart from introducing stochastic Lie systems, this paper suggests the applicability of various generalisations of Lie systems [20] to the stochastic domain. This unveils a new vast realm of potential applications in physics, mathematical, and biological models, offering promising avenues for further exploration. One example is the extension of the theory of foliated Lie systems [21] to the realm of stochastic differential equations. In fact, we suggest that, in analogy with the deterministic case, this can be the case when studying certain problems of our here devised stochastic energy-momentum method [25].
@@
By the stochastic Lie theorem, every stochastic Lie theorem admitting $\ell$ independent random variables is related to a family of $\ell$ vector fields (if the time is considered as a special ‘deterministic’ random variable) parametrised by random variables. All of them can be understood as linear combinations of elements of a finite-dimensional Lie algebra of vector fields (a so-called Vessiot–Guldberg Lie algebra) with coefficients depending on the random variables. More particularly, we study stochastic Lie systems admitting a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields relative to some compatible differential geometric structure: the Hamiltonian stochastic Lie systems. In particular, we mainly focus on Hamiltonian stochastic Lie systems relative to symplectic forms, although our theory is easy generalisable to other geometric structures and stochastic Lie systems. In this context, the coalgebra method is extended to Hamiltonian stochastic Lie systems to derive superposition rules. This provides an extension to the stochastic realm of the theory of Hamiltonian Lie systems and their generalisations [23].
@@
Our results are illustrated with many new examples of stochastic Lie systems. In particular, we study SIS models [32]. SIS models are epidemiological models that assume that individuals do not acquire immunity after infection. They concern two variables/compartments: $S$ representing susceptible individuals and $I$ representing infected individuals in a large population of size $N$ where a single disease is spreading. SIS systems are usually employed in the literature in a deterministic manner [27]. This can be used to describe some of their features, but not all, as some are of pure stochastic nature. Our models can also be used to study stochastic models appearing in Lotka–Volterra models, Palomba–Goodwin models [49], stochastic oscillators [35, 36], etc. In fact, stochastic Lie systems in general and Hamiltonian stochastic Lie systems, in particular, seem to have a quite large field of potential applications.
@@
We are also concerned with a theory of stability for Hamiltonian stochastic systems and, in particular, Hamiltonian stochastic Lie systems. Our study is specially concerned with linear ones, which appear as an approximation of nonlinear ones and they share some stability properties with them [4]. Moreover, the basis for a energy-momentum method [46] for Hamiltonian stochastic differential equations is laid down by using some results in [40]. In particular, this allows one to study the relative equilibrium points of Hamiltonian systems, i.e. points where the dynamics is generated by Hamiltonian symmetries of the system under study. A characterisation (see Theorem 6.5) of the relative equilibrium points for stochastic Hamiltonian systems in terms of critical points of its Hamiltonian functions is developed. As a byproduct, strong constants of motion [40] for Hamiltonian stochastic Lie systems are briefly studied and illustrated with examples.
@@
The structure of the paper goes as follows. Section 2 is a brief introduction to stochastic differential equations, stochastic Lie systems, and many other notions to be used in this paper. It also shows the difference of the form of stochastic Lie systems in the Stratonovich and the Itô approaches. Section 3 is concerned with superposition rules for stochastic Lie systems and it reviews and corrects the previous version of the stochastic Lie system. Section 4 deals with Hamiltonian stochastic Lie systems. In particular, we define the newly proposed stochastic foliated Lie systems, and the Hamiltonian counterparts of stochastic Lie systems. A theory of stability for stochastic Lie systems is given in Section 5. A relative equilibrium notion is presented and studied in Section 6 and a stochastic version of a classical result characterising relative equilibrium points is presented. Meanwhile, Section 7 develops the Poisson coalgebra method for Hamiltonian stochastic Lie systems. Finally, Section 8 presents our conclusions and future work.
@@ 2 Stochastic differential equations and stochastic Lie systems
This section provides a concise introduction to stochastic differential equations and stochastic Lie systems. A detailed survey on the theory of stochastic differential equations can be found in [4, 26], while the theory of the stochastic Lie systems was elaborated for the first time in [39], which offers a precise formulation of the theory. To avoid technical details, we will assume objects to be smooth, locally defined, and problems at generic points satisfying very mild conditions. Following the classical Lie systems theory [17, 20, 44], we here provide a stochastic Lie system definition not based on the superposition rule notion.
@@
for certain functions $X^{1},\dotsc,X^{n}\in\mathscr{C}^{\infty}(\mathbb{R}\times M)$, is deterministic in the sense that an initial condition in $M$ at a time $t_{0}$ establishes, under mild conditions on the functions $X^{1},\ldots,X^{n}$, a unique solution giving a position in $M$ for every $t\in\mathbb{R}$. Then, (2.1) may be modified by considering that its form can also depend on certain ‘stochastic processes’ $B^{1},\ldots,B^{r}$, namely a series of $t$-dependent random variables satisfying certain appropriate conditions to be set hereafter in detail. This fact is shown by the expression
@@
for functions $X^{i}_{1},\ldots,X_{\ell}^{i}\in\mathscr{C}^{\infty}(\mathbb{R}^{\ell}\times M)$, with $B=(t,B^{2},\ldots,B^{\ell})$, $\Gamma=(\Gamma^{1},\ldots,\Gamma^{n})$, and $i=1,\ldots,n$. The symbol $\circ$ has been used to indicate that (2.2) is understood in the so-called Stratonovich interpretation, to be briefly explained afterwords. More precisely, $(\Omega,\mathcal{F},P)$ is a probability space, where $\Omega$ is a manifold, $\mathcal{F}$ is a $\sigma$-algebra of subsets of $\Omega$, and $P:\mathcal{F}\rightarrow[0,1]$ is a probability function on $\mathcal{F}$. Each $B_{\alpha}:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$ is a semi-martingale for $\alpha=1,\ldots,r$. Semi-martingales are good integrators relative to the Itô or the Stratonovich integrals due to its properties. In short, a martingale is a sequence of stochastic processes such that, at a particular time, the conditional expectation of the next value is equal to the present value, independently of all previous values.
A stochastic differential equation is then an expression on a manifold $M$ of the form [^1] Since $\mathfrak{S}$ stands for a Stratonovich operator, we will sometimes write $\mathfrak{S}(X,\Gamma)\delta X$ instead of $\mathfrak{S}(X,\Gamma)\circ\delta X$ to simplify the notation when it is clear what we mean from context.
@@
where $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is an $\mathbb{R}^{\ell}$-valued semi-martingale and $\mathfrak{S}(B,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$, with $(B,\Gamma)\in\mathbb{R}^{\ell}\times M$, describes a Stratonovich operator. Every basis in $\mathrm{T}^{\ast}\mathbb{R}^{\ell}$ allows one to decompose $\mathfrak{S}(B,\Gamma)$ into $\ell$-components $(\mathfrak{S}_{1}(B,\Gamma),\ldots,\mathfrak{S}_{\ell}(B,\Gamma))$ in the chosen basis, which will be frequently employed hereafter. Every particular solution to (2.3) is also a semi-martingale $\Gamma:\mathbb{R}_{+}\times\Omega\rightarrow M$. Moreover, we say that a particular solution has initial condition $\Gamma_{0}\in M$ when $\Gamma(0,\omega_{0})=\Gamma_{0}$ for every $\omega_{0}\in\Omega$ with probability one. Note that the standard time can be considered as a random variable $t:(t,\omega_{0})\in\mathbb{R}_{+}\times\Omega\mapsto t\in\mathbb{R}$, whose value is independent of $\Omega$ and to be included as the first component of $B$. In practice, $B$ is related to Brownian motions, also called Wiener processes, càdlág martingales, Itô processes of the form $\delta X=\sigma\delta W+\mu\delta t$ for a Brownian motion $W$ and adapted processes $\sigma,\mu$, Lévy processes, etc. It is worth noting that semimartingales form the largest class of processes for which the Itô integral can be defined. On the other hand, white noises are not semi-martingales.

### v3

@@ Abstract
This paper provides a practical approach to stochastic Lie systems, i.e. stochastic differential equations whose general solutions can be written as a function depending only on a generic family of particular solutions and some constants related to initial conditions. We correct the stochastic Lie theorem characterising stochastic Lie systems, proving that, contrary to previous claims, it retains its classical form in the Stratonovich approach. Meanwhile, we show that the form of stochastic Lie systems may significantly differ from the classical one in the Itô formalism. New generalisations of stochastic Lie systems, like the so-called stochastic foliated Lie systems, are introduced. Subsequently, we focus on stochastic Lie systems that are Hamiltonian systems relative to different geometric structures. Special attention is paid to the symplectic case. We study their stability properties and lay the foundations of a stochastic energy-momentum method. A stochastic Poisson coalgebra method is developed to derive superposition rules for Hamiltonian stochastic Lie systems. Potential applications of our results are presented for biological stochastic models, stochastic oscillators, stochastic Lotka–Volterra systems, Palomba–Goodwin models, among others. Our findings complement previous approaches by using stochastic differential equations instead of deterministic equations designed to capture some of the features of models of stochastic nature.
Keywords: energy-momentum method, Hamiltonian stochastic Lie system, Poisson coalgebra, stability, stochastic Lie system, Stratonovich formalism, superposition rule.
@@ 1 Introduction
In its most classical definition, a superposition rule is a $t$-independent function that describes the general solution of a $t$-dependent system of first-order ordinary differential equations (ODEs) in normal form, a so-called Lie system, via a generic family of its particular solutions and a set of constants related to initial conditions [13, 18, 19, 68]. Superposition rules are used, for instance, in approximate and numerical methods, as they are applicable to Lie systems whose exact general solutions in explicit form are unknown [23, 62, 68]. In particular, the knowledge of a particular finite set of exact and/or approximate solutions of a Lie system permits one to study its general solution via superposition rules [68].
@@
The superposition rule concept traces its origins back to Sophus Lie’s pioneering and celebrated book [52], edited by Georg Scheffers. In that work, Lie stated the theorem nowadays called the Lie theorem, characterising Lie systems. Prior to that, Lie briefly described his Lie theorem in [51] without a proof, as a criticism of some previous works on superposition rules by Vessiot, Guldberg, Köningsberger, and other researchers (see [26, 40, 66] and references therein). Lie stated that the results on superposition rules for differential equations on $\mathbb{R}$ devised by previous authors were a simple application of his theory on infinitesimal transformation groups. Moreover, [52] laid down the foundations for the theory of Lie systems. The Lie theorem is also called Lie–Scheffers theorem, as Scheffers took part in editing Lie’s work [52], or Lie superposition theorem [14]. Previous remarks suggest saying ‘Lie theorem’ rather than ‘Lie–Scheffers theorem’ and using the denomination ‘Lie system’ instead of ‘Lie–Scheffers system’. Since Vessiot made important contributions to the theory of Lie systems [23, 26], the term ‘Lie–Vessiot system’ is also an appropriate designation for a Lie system. Scheffers, however, never independently researched Lie systems, and it is quite unlikely that he established any findings on the subject on his own. [26].
@@
The Lie theorem states that a $t$-dependent system of ODEs in normal form admits a superposition rule if and only if it describes the integral curves of a $t$-dependent family of vector fields that can be viewed as a curve in a finite-dimensional Lie algebra of vector fields, a Vessiot–Guldberg Lie algebra (see [14, 18, 19] for modern approaches and further details).
Lie systems have been thoroughly studied due to their widespread occurrence in physics and mathematics (see [23, 26], which contain more than 200 references on Lie systems and related topics). In the 1980s, Winternitz and his colleagues at the Centre de Recherches Mathématiques of the University of Montreal conducted an extensive study of Lie systems. In subsequent years, while Winternitz shifted his focus to other subjects, some of his collaborators continued exploring the topic [57]. Additionally, scholars from Poland, Italy, Spain, Mexico, and Russia, such as J. Grabowski, G. Marmo, J. F. Cariñena, J. de Lucas, F. J. Herranz, R. Flores-Spinoza, Y. Vorobiev and N. H. Ibragimov together with their research teams, began contributing to this field (see [26, Chapter 1], [23] and references therein).
There has been a vast effort to generalise Lie systems to more general situations: $t$-dependent Schrödinger equations [21, 23], partial differential equations [57], quasi–Lie systems [23], foliated Lie systems [18, 69], discrete differential equations [11, 61], stochastic differential equations [48], superdifferential equations [10], and others [23]. There has also been much interest in describing the geometrical properties of Lie systems and in using them to study differential geometric problems (see [3, 4, 17, 26, 32] and references therein). Moreover, Lie systems are related to important physical and mathematical models, which strongly motivates their analysis [18, 20, 23, 31, 33]. In this work, we are mainly concerned with the extension of superposition rules and Lie systems to the realm of stochastic differential equations [48]. In this way, we aim to draw the attention of researchers working on Lie systems to stochastic models, and vice versa. Therefore, to enhance accessibility to our work, we will provide a concise overview of various geometric and stochastic concepts.
@@
Stochastic differential equations may describe phenomena that deterministic differential equations cannot [2]. For example, the probability of disease extinction or outbreak, the quasi-stationary probability distribution, the final size distribution, and the expected duration of an epidemic are features that cannot effectively be modelled by deterministic methods [2]. These and other reasons motivate the great interest in studying stochastic differential equations. Nevertheless, some deterministic differential equations can capture certain interesting characteristics of models even more easily than stochastic differential equations [16], and stochastic differential equations offer a complementary alternative view [36].
@@
It is interesting to extend superposition rules to stochastic differential equations. The work [48] extends the notion of a superposition rule to a class of systems of stochastic first-order ordinary differential equations, called stochastic Lie systems. The authors use a Stratonovich approach, as this makes the theory similar to the deterministic theory of Lie systems [18, 23, 26]. Moreover, [48] presents a precise and interesting account of certain local results about stochastic Lie systems and, as a byproduct, it also explains many technical results on standard Lie systems, which are usually absent in the literature [18, 19, 26, 63]. Despite its mathematical interest, many of the technical details given in [48] are frequently omitted, as they generally have few practical applications. As noted in [14, 26], standard works on Lie systems, even theoretical ones, are essentially interested in local aspects and generic points, which leads them to skip many technical proofs analysed in [48]. Nevertheless, there is mathematical interest in both global and local technical aspects, as illustrated by [14] for global superposition rules and by [48]. It is worth stressing that stochastic Lie systems were found to have applications in the description of Brownian motions, economic models such as the Black–Scholes theory of derivative pricing, and so on [48]. The potential interest of stochastic Lie systems in epidemic models was very briefly mentioned in [16], without, as far as we know, any further development. It is worth noting that there are many new potential applications of the theory of Lie systems to stochastic models, which have so far remained almost unexplored.
@@
In [48], a stochastic Lie theorem characterising Stratonovich stochastic first-order ordinary differential equations admitting a superposition rule was devised, but it contains a mistake that changes its meaning and potential applications. More precisely, the direct part of the stochastic Lie theorem in [48] states that a stochastic Lie system may admit, locally, a superposition rule if its associated Stratonovich operator is related to a family of vector fields spanning an involutive distribution. Our present work proves that the vector fields must additionally span a finite-dimensional real Lie algebra of vector fields, which is a much stronger condition that already appears in the classical Lie–Scheffers theorem [18, 19, 52]. We also determine the precise point of the mistake in [48], correct the statement of the stochastic Lie theorem, and provide a counterexample, based on SIS epidemic models, to illustrate when the stochastic Lie theorem in [48] fails. It is worth stressing that the differences between the incorrect and the correct versions of the stochastic Lie theorem have no impact on the applications carefully studied in [48].
@@
Moreover, our work presents a concise introduction to stochastic Lie systems, aiming to provide a practical approach while avoiding technical details that are not necessary for general purposes. In this sense, it simplifies the elegant and mostly rigorous mathematical treatment in [48] by using standard assumptions in mathematical constructions. For instance, we focus on local results at generic points, which significantly simplifies previously required techniques.
@@
Our correction of the stochastic Lie theorem shows that the stochastic Lie theorem has no exclusive stochastic features in the Stratonovich approach: it retains the conditions of the classical Lie theorem. Meanwhile, it should be stressed that the conditions for a stochastic differential equation in Itô form [65] to become a stochastic Lie system do not follow the standard form expected from the deterministic Lie theorem [18, 19, 23]. This is very important in practice, as many relevant stochastic differential equations are given in Itô form and must be translated into the Stratonovich approach [64] in order to apply the methods of our work and of [48]. In this respect, the relation between the Itô and the Stratonovich approaches is reviewed, and some examples with potential applications are provided in this work. It is worth noting that stochastic differential equations in Itô form may appear to be stochastic Lie systems, but they are not. This occurs in certain SIS epidemiological models [35], as shown in this work.
@@
Apart from introducing stochastic Lie systems, this paper suggests the applicability of various generalisations of Lie systems [23] to the stochastic domain. This opens a new vast realm of potential applications in physical, mathematical, and biological models, offering promising avenues for further exploration. One example is the extension of the theory of foliated Lie systems [69] to the realm of stochastic differential equations. In fact, we suggest that, in analogy with the deterministic case, this may arise when studying certain problems of our stochastic energy-momentum method devised here [54].
@@
By the stochastic Lie theorem, every stochastic Lie system admitting $\ell$ independent random variables is related to a family of $\ell$ vector fields (if time is considered as a special ‘deterministic’ random variable) parametrised by random variables. All of them can be understood as linear combinations of elements of a finite-dimensional Lie algebra of vector fields (a so-called Vessiot–Guldberg Lie algebra) with coefficients depending on the random variables. More particularly, we study stochastic Lie systems admitting a Vessiot–Guldberg Lie algebra of Hamiltonian vector fields relative to some compatible differential geometric structure: the Hamiltonian stochastic Lie systems. In particular, we mainly focus on Hamiltonian stochastic Lie systems relative to symplectic forms, although our theory is easily generalisable to other geometric structures and stochastic Lie systems. In this context, the coalgebra method is extended to Hamiltonian stochastic Lie systems to derive superposition rules. This provides an extension to the stochastic realm of the theory of Hamiltonian Lie systems and their generalisations [26].
@@
Our results are illustrated with many new examples of stochastic Lie systems. In particular, we study SIS models [35]. SIS models are epidemiological models that assume that individuals do not acquire immunity after infection. They concern two variables/compartments: $S$, representing susceptible individuals, and $I$, representing infected individuals in a large population of size $N$ where a single disease is spreading. SIS systems are usually treated in the literature in a deterministic manner [16]. This can be used to describe some of their features, but not all, as some are purely stochastic in nature. Our models can also be used to study stochastic models arising in Lotka–Volterra systems, Palomba–Goodwin models [60], stochastic oscillators [41, 42], etc. In fact, stochastic Lie systems in general, and Hamiltonian stochastic Lie systems in particular, seem to have a wide range of potential applications.
@@
We are also concerned with a theory of stability for Hamiltonian stochastic systems and, in particular, Hamiltonian stochastic Lie systems. Our study is specially concerned with linear ones, which appear as approximations of nonlinear ones and share some stability properties with them [7]. An example of a stochastic oscillator with a drift is analysed. Moreover, the basis for an energy-momentum method [54] for Hamiltonian stochastic differential equations is established by using some results in [46]. In particular, this allows one to study the relative equilibrium points of Hamiltonian systems, i.e. points where the dynamics is generated by Hamiltonian symmetries of the system under study. A characterisation (see Theorem 6.5) of the relative equilibrium points for stochastic Hamiltonian systems in terms of critical points of their Hamiltonian functions is developed. As a byproduct, strong constants of motion [46] for Hamiltonian stochastic Lie systems are briefly studied and illustrated with examples.
@@
The structure of the paper is as follows. Section 2 is a brief introduction to stochastic differential equations, stochastic Lie systems, and many other notions to be used in this paper. It also shows the difference between the form of stochastic Lie systems in the Stratonovich and the Itô approaches. Section 3 is concerned with superposition rules for stochastic Lie systems and reviews and corrects the previous version of the stochastic Lie theorem. Section 4 deals with Hamiltonian stochastic Lie systems. In particular, we define the newly proposed stochastic foliated Lie systems and the Hamiltonian counterparts of stochastic Lie systems. A theory of stability for stochastic Lie systems is given in Section 5. A relative equilibrium notion is presented and studied in Section 6, and a stochastic version of a classical result characterising relative equilibrium points is presented. Meanwhile, Section 7 develops the Poisson coalgebra method for Hamiltonian stochastic Lie systems. Finally, Section 8 presents our conclusions and future work.
@@ 2 Stochastic differential equations and stochastic Lie systems
This section provides a concise introduction to stochastic differential equations and stochastic Lie systems. We have tried to provide enough information to follow the paper for people working on stochastic differential equations or the theory of Lie systems.
A detailed survey on the theory of stochastic differential equations can be found in [7, 28], while the theory of stochastic Lie systems was elaborated for the first time in [48], which offers a precise formulation of the theory. To avoid technical details, we will assume objects to be smooth, locally defined, and problems at generic points satisfying very mild conditions. Following the classical Lie systems theory in [23, 26], we here provide a definition of a stochastic Lie system not based on the notion of a superposition rule.
@@
for certain functions $X^{1},\dotsc,X^{n}\in\mathscr{C}^{\infty}(\mathbb{R}\times M)$, is deterministic in the sense that an initial condition in $M$ at a time $t_{0}$ establishes, under mild conditions on the functions $X^{1},\ldots,X^{n}$, a unique solution giving a position in $M$ for every $t\in\mathbb{R}$. Geometrically, the coefficients $X^{1}(t,\Gamma),\ldots,X^{n}(t,\Gamma)$ give rise to a $t$-parametrised vector field on $M$ of the form
$X=\sum_{i=1}^{n}X^{i}(t,\Gamma)\frac{\partial}{\partial\Gamma^{i}},$
which is formally called a $t$-dependent vector field. More generally, an $\mathbb{R}^{\ell}$-dependent vector field on a manifold $M$ is a family of vector fields on $M$ parametrised by elements of $\mathbb{R}^{\ell}$ (see [23] for details).
Then, (2.1) may be modified by considering that its form can also depend on certain ‘stochastic processes’ $B^{1},\ldots,B^{r}$, namely a series of $t$-dependent random variables satisfying certain appropriate conditions to be set hereafter in detail. This fact is shown by the expression
@@
for functions $X^{i}_{1},\ldots,X_{\ell}^{i}\in\mathscr{C}^{\infty}(\mathbb{R}^{\ell}\times M)$, with $B=(t,B^{2},\ldots,B^{\ell})$, $\Gamma=(\Gamma^{1},\ldots,\Gamma^{n})$, and $i=1,\ldots,n$. The symbol $\circ$ has been used to indicate that (2.2) is understood in the so-called Stratonovich interpretation, to be briefly explained afterwards. More precisely, $(\Omega,\mathcal{F},P)$ is a probability space, where $\Omega$ is a manifold, $\mathcal{F}$ is a $\sigma$-algebra of subsets of $\Omega$, and $P:\mathcal{F}\rightarrow[0,1]$ is a probability function on $\mathcal{F}$. Each $B_{\alpha}:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}$ is a semi-martingale for $\alpha=1,\ldots,r$. Semi-martingales are good integrators relative to the Itô or the Stratonovich integrals due to their properties. In short, a martingale is a sequence of stochastic processes such that, at a particular time, the conditional expectation of the next value is equal to the present value, independently of all previous values.
A stochastic differential equation is then an expression on a manifold $M$ of the form
@@
where $B:\mathbb{R}_{+}\times\Omega\rightarrow\mathbb{R}^{\ell}$ is an $\mathbb{R}^{\ell}$-valued semi-martingale and $\mathfrak{S}(B,\Gamma):\mathrm{T}_{B}\mathbb{R}^{\ell}\to\mathrm{T}_{\Gamma}M$, with $(B,\Gamma)\in\mathbb{R}^{\ell}\times M$, describes a Stratonovich operator. Every basis in $\mathrm{T}^{\ast}\mathbb{R}^{\ell}$ allows one to decompose $\mathfrak{S}(B,\Gamma)$ into $\ell$-components $(\mathfrak{S}_{1}(B,\Gamma),\ldots,\mathfrak{S}_{\ell}(B,\Gamma))$ in the chosen basis, which will be frequently employed hereafter. Geometrically, every component is a mapping
$\mathfrak{S}_{\alpha}:(B,\Gamma)\in\mathbb{R}^{\ell}\times M\mapsto\mathfrak{S}_{\alpha}(B,\Gamma)\in\mathrm{T}_{\Gamma}M\subset\mathrm{T}M,\qquad\alpha=1,\ldots,\ell,$
which gives rise, for every fixed value of $B$, to a vector field on $M$. In fact, $\mathfrak{S}_{\alpha}(B,\cdot):\Gamma\in M\mapsto\mathfrak{S}_{\alpha}(B,\Gamma)\in\mathrm{T}_{\Gamma}M\subset\mathrm{T}M$. In other words, $\mathfrak{S}_{\alpha}$ can be understood as an $\mathbb{R}^{\ell}$-parametrised vector field, which is usually called an $\mathbb{R}^{\ell}$-dependent vector field [23]. Every particular solution to (2.3) is also a semi-martingale $\Gamma:\mathbb{R}_{+}\times\Omega\rightarrow M$. Moreover, we say that a particular solution has initial condition $\Gamma_{0}\in M$ when $\Gamma(0,\omega_{0})=\Gamma_{0}$ for every $\omega_{0}\in\Omega$ with probability one. Note that the standard time can be considered as a random variable $t:(t,\omega_{0})\in\mathbb{R}_{+}\times\Omega\mapsto t\in\mathbb{R}$, whose value is independent of $\Omega$ and is included as the first component of $B$. In practice, $B$ is related to Brownian motions, also called Wiener processes, càdlág martingales, Itô processes of the form $\delta X=\sigma\delta W+\mu\delta t$ for a Brownian motion $W$ and adapted processes $\sigma,\mu$, Lévy processes, etc. It is worth noting that semi-martingales form the largest class of processes for which the Itô integral can be defined. On the other hand, white noises are not semi-martingales.
We assume that the driving processes are adapted to the natural filtration $(\mathcal{F}_{t})_{t\geq 0}$, i.e. we only allow processes that evolve consistently with the information available up to the present time. Vector fields will satisfy local Lipschitz and linear growth conditions (or equivalent geometric hypotheses) guaranteeing existence and uniqueness of solutions. These conditions are a stochastic analogue of the usual ODE assumptions, ensuring existence and uniqueness of solutions to the SDE. Unless otherwise stated, the semi-martingales considered have continuous trajectories (in particular, Brownian motions). If jump processes such as Lévy noise are allowed, the canonical Stratonovich calculus is not well defined; in that case, the Marcus integral provides the appropriate extension. Since our results rely on the Stratonovich calculus on smooth manifolds, we restrict throughout to continuous semi-martingale drivers.